Opened 10 years ago

Closed 9 years ago

#214 closed flight processing (fixed)

VOC_11, flight day 315/2008, Chile

Reported by: benj Owned by: mggr
Priority: alpha 4 high Milestone: 2008 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description

Data location: /data/larsenb1/scratch/arsf/vocals/vocals-Q/voc11-31508 (symlink in ~arsf/arsf_data/2008/flight_data/vocals/VOC_11-2008_315_Chile)

Data arrived from ARSF via SATA disk Q on 12th December 2008.

Scientific details: VOCALS cloud campaign

Priority:

PI: G. Vaughan

Sensors:

  • Eagle
  • Hawk

Quicklook

TBD.

Change History (15)

comment:1 Changed 10 years ago by jatt

Have renamed Eagle files so all start with VNIR prefix.

comment:2 Changed 10 years ago by jatt

  • Owner changed from jatt to peland

Delivery directory created at ~arsf/arsf_data/2008/flight_data/vocals/VOC_11-2008_315_Chile/delivery/20090518/VOC11

comment:3 Changed 10 years ago by jatt

All flight lines are CM 69.

The pitch, roll and heading values used for the flight are:-

Sensor: Eagle

Flight Line Pitch Roll Heading
01-11 -0.2 -0.45 -0.5

Sensor: Hawk

Flight Line Pitch Roll Heading
01 -0.21 0.13 0.767
02 -0.287 0.294 0.821
03 -0.201 0.337 0.746
04 -0.206 0.334 0.771
05 -0.274 0.334 0.765
06 -0.238 0.334 0.757
07 -0.121 0.339 0.849
08 -0.257 0.159 0.795
09 -0.254 0.17 0.785
10 -0.277 0.121 0.618
11 -0.245 0.134 0.816

comment:4 Changed 10 years ago by peland

Logfile OK, all L1 files normal.

Tried azgcorr on e315011b.hdf, terminated due to turn in flightline. e315021b.hdf seems to run normally and gives sensible output.

Ticket ready to close?

comment:5 Changed 10 years ago by benj

  • Owner changed from peland to jatt

Leave the ticket for now (we tend to hand them off to Mike to go to NEODC and then he closes them when that's done). Normally we rsync the working copy back to the repository but for VOCALS we've been working directly in the repository due to lack of disk space. Passing the ticket back to Jane for delivery.

comment:6 Changed 10 years ago by jatt

All data delivered 9th June 2009

comment:7 Changed 10 years ago by mggr

  • Type changed from task to flight processing

comment:8 Changed 10 years ago by mggr

  • Component changed from Processing: general to Archiving
  • Owner changed from jatt to mggr

Archived locally by DTG on 29/Jan/2009. Ready to go to NEODC.

comment:9 Changed 10 years ago by benj

Actually archived on 15th June 1009. Still OK to go to NEODC.

comment:10 Changed 9 years ago by emca

Data re-delivered per request from L. Labrador 22/12/09

comment:11 Changed 9 years ago by benj

Copied to SATA disk "Linex backup 1" (sic) 19/01/2009, ready for delivery to NEODC

comment:12 Changed 9 years ago by benj

Disregard previous comment, not actually on that disk (disk space problems).

comment:13 Changed 9 years ago by benj

Copied to SATA disk H 27/01/2009, ready for delivery to NEODC

comment:14 Changed 9 years ago by benj

Disk posted to NEODC 18/02/2010

comment:15 Changed 9 years ago by benj

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.